one number is 7 more than another number. Find the two numbers if three times the larger number exceeds four times the smaller number by 5.
Answer by checkley77(12844)

X=Y+7
3X>4Y+5
3(Y+7)>4Y+5
3Y+21>4Y+5
3Y-4Y>5-21
-Y>-16
Y<16 {NOTE: THE SIGN CHANGE WHEN DIVIDING BY A NEGATIVE VALUE.
X=16+7=23
PROOF:
LET Y=15
3*23>4*15+5
69>60+5
69>65